A line passes through the points (7,-4,8) and (9,0,5)

i) Determine the parametric and symmetric equations of the line

ii)Find any other point on the line

iii)Show that the line does not pass through the origin



1
Expert's answer
2021-08-16T15:34:49-0400

i) The collinear vector of a line is "\overline{v}=(9-7,0-(-4),5-8)=(2,4,-3)." Then the symmetric equation of the line is

"\frac{x-7}{2}=\frac{y+4}{4}=\frac{z-8}{-3}."

It follows that the parametric equation of the line is

"\begin{cases}x=7+2t\\y=-4+4t\\z=8-3t\end{cases}."


ii) For "t=-1" we have that

"x=7+2(-1)=5,\ y=-4+4(-1)=-8,\ z=8-3(-1)=11."

Then the point "(5,-8,11)" is on the line.


iii) If "y=0," then "-4+4t=0," and hence "t=1." On the other hand for "t=1," we have that "x=9\ne 0," and hence the line does not pass through the origin.